An aqueous solution of acetic acid is found to have the following concentrations at 25°C: [CH3COOH] = 1.65 x 10-2 M; [H30+] = 5.44 x 10-4 M; and [CH3COO-] = 5.44 x 10+ M. Calculate the equilibrium at 25°C. The reaction is:


CH3COOH(aq) + H2O(I) ⇆ H3O(aq) + CH3COO-(aq)



Strategy:

Using the balanced chemical equation, write the equilibrium constant expression, Kc, then substitute the given equilibrium concentrations to it.


1
Expert's answer
2022-06-06T09:25:03-0400

CH3COOH + H2O = H3O+ + CH3COO-


"K_c = \cfrac {[H_3O^+][CH_3COO^-]}{[CH_3COOH]}"



"K_c = \frac {[5,44*10^{-4}][5,44*10^{-4}]}{[1,65*10^{-2}]}= 1,8*10^{-5}"



ANSWER: 1,8 * 10-5